two cars leave a intersection at the same time. one is headed south at a constant speed of 30 mph, the other is headed west at a constant speed of 40 mph. what is the distance between the cars in 30 minutes?
Answer by checkley77(12844)

a^2+b^2=c^2
(30*.5)^2+(40*.5)^2=c^2
15^2+20^2=c^2
225+400=c^2
c^2=625
c=sqrt625
c=25 miles apart.